# Spectral Logit Sculpting: Adaptive Low-Rank Logit Transformation for Controlled Text Generation

### v1

@@ Abstract
footnotetext: *Equal contribution.Corresponding author (mhan}@zju.edu.cn).
© 2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) on complex tasks has attracted significant attention. They have demonstrated strong capabilities in mathematical reasoning [1], code generation [2], solving scientific problems [3] and supporting agent-based applications [19], especially after being fine-tuned on domain-specific data. Recently, entropy-based information has been widely applied in the field of LLM fine-tuning, serving as an effective means to enhance model performance. For instance, the step entropy compression method guides model fine-tuning by replacing low-entropy reasoning steps with special tokens, thereby compressing the reasoning chain. This approach significantly improves inference efficiency while maintaining high accuracy [5]. Additionally, the one-shot entropy minimization method only requires one unlabeled sample to update the parameters through ten steps of entropy minimization, and it can achieve performance comparable to that of thousands of steps of reinforcement learning [4].
@@ 2.2 Entropy-Based Activation of Spectral Estimation
The leading $m$ right singular vectors, denoted as $\mathbf{U}_{b}^{*}=\mathbf{V}_{b}(:,1:m)\in\mathbb{R}^{K\times m}$, capture the principal directions of variation in the recent logit history. These vectors form an orthonormal basis that spans the subspace of most significant logit variations, providing a low-rank representation essential for subsequent logit adjustment. This dimensionality reduction not only enhances computational efficiency but also focuses the adjustment on the most relevant directions of variation.
@@ 2.3 Adaptive Logit Rescaling and Reconstruction
where $\gamma\in(0,1]$ is a damping factor for the residual component. The adjusted logits $\hat{\mathbf{z}}_{t}$ replace the original top-$K$ values before sampling, effectively reshaping the distribution toward the dominant spectral directions only in high-uncertainty contexts. This approach selectively enhances consistency with recent patterns when the model shows uncertainty, while preserving original predictions when the model is already confident.
@@ 4.1 Comparative Evaluation of SLS
Importantly, SLS also attains the best performance (24.2%) on the UGPhysics science benchmark. For the Qwen2.5-7B-Instruct model, SLS is the only method that improves performance (24.2%) compared to the baseline (23.4%), while all other methods result in degradation. For the mathematically specialized Qwen2.5-Math-7B-Instruct, SLS achieves the smallest performance decline (18.3%) among all compared methods. The superiority of SLS is further confirmed on the mathematically tuned Qwen2.5-Math-7B-Instruct model. SLS achieves the best average math score (45.4%), with top performance on AMC (55.4%) and AIME (12.2%). This confirms that SLS provides complementary benefits even to specialized models by further optimizing their inference-time decoding process.

### v2

@@ Abstract
footnotetext: © 2026 IEEE. Published in ICASSP 2026 – 2026 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), scheduled for 3–8 May 2026 in Barcelona, Spain. Personal use of this material is permitted. However, permission to reprint/republish this material for advertising or promotional purposes or for creating new collective works for resale or redistribution to servers or lists, or to reuse any copyrighted component of this work in other works, must be obtained from the IEEE. Contact: Manager, Copyrights and Permissions / IEEE Service Center / 445 Hoes Lane / P.O. Box 1331 / Piscataway, NJ 08855-1331, USA. Telephone: +1 908 562 3966.
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) on complex tasks has attracted significant attention [23, 17, 21, 16, 28]. They have demonstrated strong capabilities in mathematical reasoning [4], code generation [12], solving scientific problems [13] and supporting agent-based applications [8, 24, 25, 20, 19, 22, 27], especially after being fine-tuned on domain-specific data. Recently, entropy-based information has been widely applied in the field of LLM fine-tuning, serving as an effective means to enhance model performance. For instance, the step entropy compression method guides model fine-tuning by replacing low-entropy reasoning steps with special tokens, thereby compressing the reasoning chain. This approach significantly improves inference efficiency while maintaining high accuracy [11]. Additionally, the one-shot entropy minimization method only requires one unlabeled sample to update the parameters through ten steps of entropy minimization, and it can achieve performance comparable to that of thousands of steps of reinforcement learning [3].
@@ 2.2 Entropy-Based Activation of Spectral Estimation
The leading $m$ right singular vectors, denoted as $\mathbf{U}_{b}^{*}=\mathbf{V}_{b}(:,1:m)\in\mathbb{R}^{K\times m}$, capture the principal directions of variation in the recent logit history. These vectors form an orthonormal basis that spans the subspace of most significant logit variations, providing a low-rank representation essential for subsequent logit adjustment. This dimensionality reduction not only enhances computational efficiency but also focuses the adjustment on the relevant directions of variation.
@@ 2.3 Adaptive Logit Rescaling and Reconstruction
where $\gamma\in(0,1]$ is a damping factor for the residual component. The adjusted logits $\hat{\mathbf{z}}_{t}$ replace the original top-$K$ values before sampling, effectively reshaping the distribution toward the dominant spectral directions only in high-uncertainty contexts. This approach selectively enhances consistency with recent patterns when the model shows uncertainty, while preserving original predictions when the model is confident.
@@ 4.1 Comparative Evaluation of SLS
Importantly, SLS attains the best performance (24.2%) on the UGPhysics science benchmark. For the Qwen2.5-7B-Instruct model, SLS is the only method that improves performance (24.2%) compared to the baseline (23.4%), while all other methods result in degradation. For the mathematically specialized Qwen2.5-Math-7B-Instruct, SLS achieves the smallest decline (18.3%) among all compared methods. The superiority of SLS is further confirmed on the mathematically tuned Qwen2.5-Math-7B-Instruct model. SLS achieves the best average math score (45.4%), with top performance on AMC (55.4%) and AIME (12.2%). This confirms that SLS provides complementary benefits even to specialized models by optimizing their inference-time decoding process.
@@ 4.2 Ablation Studies
4.3 Efficiency Analysis
We evaluate efficiency on Math500 with Qwen2.5-7B-Instruct. Despite generating a similar number of tokens to EM-INF ($\sim$296K vs. $\sim$306K), SLS achieves higher inference throughput (17.72 vs. 16.81 tokens/s) and computational throughput (38.0 vs. 35.6 GFLOPS). This gain stems from SLS’s lightweight design, where spectral updates are conditionally triggered only when entropy exceeds a threshold, avoiding constant computational overhead.